OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (mathstudent55):

\( 3 \sqrt{2x} + x \sqrt{8x} - 5 \sqrt{18x} \) \( = 3 \sqrt{2x} + x \sqrt{4 \times 2x} - 5 \sqrt{9 \times 2x} \) \(= 3 \sqrt{2x} + 2x \sqrt{2x} - 5 \times 3\sqrt{2x} \) \(= 3 \sqrt{2x} + 2x \sqrt{2x} - 15 \sqrt{2x} \) \(= (3 + 2x - 15)\sqrt{2x} \)

OpenStudy (anonymous):

So which step would nakim have messed up on?

OpenStudy (mathstudent55):

What Nakim did wrong was that he went from: \( (3 + 2x - 15) \sqrt{2x} \) to: \(-10x \sqrt{2x} \) \( (3 + 2x - 15) \sqrt{2x} \) \(= (2x - 12) \sqrt{2x} \) \(3 + 2x - 15 = 2x - 12 \) \(3 + 2x - 15 \ne -10x \) You can only combine like terms, and \(2x\) and \(-12x\) are not like terms.